A 1500 m3 ballon weights 1650 kg with all its equipment. It is filled with helium of density 0.2 kg m-3. If the density of air is 1.3 kg m-3, then the pull on the rope tied to the ballon will be ?

Grade:12th pass

3 Answers

Arun
4 years ago
 A floating object displaces its own weight/mass of the fluid it floats in.
air density = 1.3 kg/m³
volume of balloon = 1500 m³
mass of balloon + mass of gas in it = 1650 kg
volume of air displaced by balloon and gas in it = (1650)/1.3 = 1269 m³
excess volume = 1500 - 1269 = 231 m³
mass of air in excess volume = 231(1.3) = 300 kg
force of gravity on excess mass of air = 300(9.81) = 2946 N 
HOPE IT HELPS
 
 
Khimraj
4 years ago
Volume of baloon: V= 1500 cubic meter densiy of air: d = 1.3 kg/cu m mass of baloon: m = 1650 kg   Buoyant force on baloon = mass of air displaced = F = V x d x g = 1500x1.3x10=19500 N thus force on baloon due to the tension in rope: 19500-1650(10) = 3000N
aswanth nayak
4 years ago
A floating object displaces its own weight/mass of the fluid it floats in.
air density = 1.3 kg/m³
volume of balloon = 1500 m³
mass of balloon + mass of gas in it = 1650 kg
volume of air displaced by balloon and gas in it = (1650)/1.3 = 1269 m³
excess volume = 1500 - 1269 = 231 m³
mass of air in excess volume = 231(1.3) = 300 kg
force of gravity on excess mass of air = 300(9.81) = 2946 N 
 
second method:
Volume of baloon: V= 1500 cubic meter densiy of air: d = 1.3 kg/cu m mass of baloon: m = 1650 kg   Buoyant force on baloon = mass of air displaced = F = V x d x g = 1500x1.3x10=19500 N thus force on baloon due to the tension in rope: 19500-1650(10) = 3000N